OpenStudy (anonymous):

A rectangular garden is 16 feet long. If you walk diagonally across the garden, you would walk 20 feet. How many feet wide is the garden?

OpenStudy (anonymous):



OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

formula a^2 +b^2 =c^2?

OpenStudy (anonymous):

yep, solve for x

OpenStudy (anonymous):

ok give me a sec

OpenStudy (anonymous):

656 ft?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

16^2 + x^2 = 20^2

OpenStudy (anonymous):

did you get that?

OpenStudy (anonymous):

yeah i got that part

OpenStudy (anonymous):

ok now just simplify a little bit

OpenStudy (anonymous):

256+x^2=400

OpenStudy (anonymous):

x^2=400-256

OpenStudy (anonymous):

x^2=144

OpenStudy (anonymous):

12ft

OpenStudy (anonymous):

and what is the last step?

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

yay ok got it! thanks

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

i need help with this one

OpenStudy (anonymous):

Carpet sells for $4 per square foot and will cost you $836 to recarpet your room. If your room is 19 feet long, how many feet wide is it?

OpenStudy (anonymous):

i got 12 ft